# SkillWrapper: Generative Predicate Invention for Skill Abstraction

### v3

@@ Appendix D Proofs
[…]
For any $\mathcal{M}$$\in{\mathcal{H}}$, define
$d_{\mathrm{compl}}({\mathcal{M}},{\mathcal{M}}^{*})\;=\;\Pr_{(s,\omega,s^{\prime})\sim\mu}\bigl[(s\in\widehat{I}_{\omega}\triangle I_{\omega})\vee(s^{\prime}\in\widehat{\beta}_{\omega}\triangle\beta_{\omega})\bigr].$
Let $n$ i.i.d. samples $\{(s_{i},\omega_{i},s^{\prime}_{i})\}_{i=1}^{n}$ be drawn from $\mu$. Then, for every $\epsilon>0$,
$\Pr\!\Big[d_{\mathrm{compl}}\bigl(\widehat{{\mathcal{M}}}_{n},{\mathcal{M}}^{*}\bigr)\leq\epsilon\Big]\;\geq\;1-|{\mathcal{H}}|\,e^{-n\epsilon},$
i.e., with high probability, $\widehat{{\mathcal{M}}}_{n}$ misses fewer than an $\epsilon$-fraction of feasible transitions under $\mu$.
Proof sketch.
For $\mathcal{M}$$\in{\mathcal{H}}$, define the true error
$\mathrm{Err}({\mathcal{M}})=d_{\mathrm{compl}}({\mathcal{M}},{\mathcal{M}}^{*}),$
and the empirical error
$\widehat{\mathrm{Err}}({\mathcal{M}})=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\!\left[(s_{i}\in\widehat{I}_{\omega}\triangle I_{\omega})\vee(s^{\prime}_{i}\in\widehat{\beta}_{\omega}\triangle\beta_{\omega})\right].$
By Lemma 2, $\widehat{\mathrm{Err}}(\widehat{{\mathcal{M}}}_{n})=0$. Suppose some $\mathcal{M}$$\in{\mathcal{H}}$ satisfies $\mathrm{Err}({\mathcal{M}})\geq\epsilon$. Then each sample has probability at least $\epsilon$ of revealing an error. The chance of seeing none in $n$ i.i.d. draws is at most $e^{-n\epsilon}$ (by Hoeffding/Chernoff).
Applying a union bound over all $\mathcal{M}$$\in{\mathcal{H}}$,
$\Pr\bigl[\exists{\mathcal{M}}\in{\mathcal{H}}:\mathrm{Err}({\mathcal{M}})\geq\epsilon\ \wedge\ \widehat{\mathrm{Err}}({\mathcal{M}})=0\bigr]\;\leq\;|{\mathcal{H}}|\,e^{-n\epsilon}.$
Since $\widehat{{\mathcal{M}}}_{n}$ has $\widehat{\mathrm{Err}}=0$, the event $\mathrm{Err}(\widehat{{\mathcal{M}}}_{n})\geq\epsilon$ is contained in this bound. Thus, with probability at least $1-|{\mathcal{H}}|e^{-n\epsilon}$, $d_{\mathrm{compl}}(\widehat{{\mathcal{M}}}_{n},{\mathcal{M}}^{*})\leq\epsilon$. ∎
Theorem 4 (Soundness of SkillWrapper).
Let $\mathcal{T}$} be the set of transition tuples $\langle s,\omega,s^{\prime}\rangle$ with $s\in{\mathcal{S}}$ and $\omega\in\Omega$. Let $B_{n}$ be an experience buffer of $n$ samples drawn from $\mathcal{T}$}, each labeled with outcome. Suppose SkillWrapper learns a model
${\mathcal{M}}_{n}=({\mathcal{P}}_{n},{\mathcal{A}}_{n}),$
where each $a\in{\mathcal{A}}_{n}$ corresponds to some skill $\omega\in\Omega$. Then, for every operator $a\in{\mathcal{A}}_{n}$ associated with $\omega$, there exists a real transition $\langle s,\omega,s^{\prime}\rangle\in B_{n}$ such that
$\underline{s}\models\textsc{Pre}_{a}\quad\text{and}\quad\underline{s}^{\prime}\models\textsc{Eff}_{a}.$
Proof sketch.
By construction, SkillWrapper derives each operator $a\in{\mathcal{A}}_{n}$ from transitions in $B_{n}$ through its wrapper procedure. If $a$ is associated with skill $\omega$, then its preconditions $\textsc{Pre}_{a}$ are obtained from the abstract representation of some observed $s$, and its effects $\textsc{Eff}_{a}$ from the corresponding $s^{\prime}$. Hence there must exist $\langle s,\omega,s^{\prime}\rangle\in B_{n}$ such that $\underline{s}\models\textsc{Pre}_{a}$ and $\underline{s}^{\prime}\models\textsc{Eff}_{a}$.
If no such transition existed, then $a$ would be unsupported by data and would not have been generated. Thus every operator in ${\mathcal{A}}_{n}$ corresponds to at least one valid observed transition, proving soundness. ∎
Proof (by Contradiction).
Assume, for contradiction, that there is an operator $a\in\mathcal{A}_{n}$ for which no transition $\langle s,\omega,s^{\prime}\rangle\in B_{n}$ supports it. That would mean:
$\text{(1) }a\text{ is in the learned model},\quad\quad$
[…]
However, SkillWrapper introduces or refines operators only in response to observed transitions $\langle s,\omega,s^{\prime}\rangle$ that cannot be explained by any existing operator in $\mathcal{A}_{n}$. Therefore, if $a$ exists in the final model, it must have been created when the system encountered a transition $\langle s,\omega,s^{\prime}\rangle$ with $\underline{s}$ and $\underline{s}^{\prime}$ not accounted for by any previously existing operator. That transition becomes the “anchor” for $a$’s preconditions and effects.
Hence, there must be at least one real transition $\langle s,\omega,s^{\prime}\rangle\in B_{n}$ matching the preconditions and effects of $a$, contradicting assumption (2). Consequently, our assumption is false, and each operator indeed has a supporting transition in $B_{n}$. This completes the proof. ∎
D.1 Illustration of Predicate Invention in Low-level Space
Following the last section, we illustrate how SkillWrapper invents new predicates under all circumstances with guarantees using the condition. We discuss the case of precondition here, and guarantees of effect follow the same logic.
$\mathbf{U}$$\mathbf{I_{\omega}}$$s_{1}$$s_{2}$
By assumption, there exist two transitions: $\langle s_{1},\omega,s_{1}^{\prime}\rangle,\langle s_{2},\omega,s_{2}^{\prime}\rangle\text{ such that }s_{1}\in I_{\omega},s_{2}\notin I_{\omega}$ initially.
There are three possible circumstances of the resulting learning model from $\langle s_{1},\omega,s_{1}^{\prime}\rangle$ and $\langle s_{2},w,s_{2}^{\prime}\rangle$. For each of them, we discuss all possible cases of more transitions with the starting state falling into each section of the state space.
(1). The learned model has $I_{\omega}\subset\alpha$. Then, for each section that initial states of new transitions can fall in:
$\mathbf{I_{\omega}}$$\mathbf{U}$$s_{1}$$s_{2}$aabbcc
  • (a): $a\notin\alpha,a\notin I_{\omega}$.
    Thus, $\nexists\>\langle s,\omega,s^{\prime}\rangle\text{ such that }\mathbf{1}_{\alpha}(a)=\mathbf{1}_{\alpha}(s)$ while $I_{\omega}(a)\neq I_{\omega}(s)$. No additional predicate need to be invented.
  • (b): $b\in\alpha,a\notin I_{\omega}$.
    Thus, $\mathbf{1}_{\alpha}(b)=\mathbf{1}_{\alpha}(s_{1})$ while $I_{\omega}(a)\neq I_{\omega}(s)$. New predicate will be invented.
  • (c): $a\in\alpha,a\in I_{\omega}$.
    Thus, $\nexists\>\langle s,\omega,s^{\prime}\rangle\text{ such that }\mathbf{1}_{\alpha}(a)=\mathbf{1}_{\alpha}(s)$ while $I_{\omega}(a)\neq I_{\omega}(s)$. No additional predicate need to be invented.
(2). The learned model has $\alpha\subset I_{\omega}$. Then, for each section that initial states of new transitions can fall in:
$\mathbf{I_{\omega}}$$\mathbf{U}$$s_{1}$$s_{2}$aabbcc
  • (a): $a\notin\alpha,a\notin I_{\omega}$.
    Thus, $\nexists\>\langle s,\omega,s^{\prime}\rangle\text{ such that }\mathbf{1}_{\alpha}(a)=\mathbf{1}_{\alpha}(s)$ while $I_{\omega}(a)\neq I_{\omega}(s)$. No additional predicate need to be invented.
  • (b): $b\notin\alpha,a\in I_{\omega}$.
    Thus, $\mathbf{1}_{\alpha}(b)=\mathbf{1}_{\alpha}(s_{1})$ while $I_{\omega}(a)\neq I_{\omega}(s)$. New predicate will be invented.
  • (c): $a\in\alpha,a\in I_{\omega}$.
    Thus, $\nexists\>\langle s,\omega,s^{\prime}\rangle\text{ such that }\mathbf{1}_{\alpha}(a)=\mathbf{1}_{\alpha}(s)$ while $I_{\omega}(a)\neq I_{\omega}(s)$. No additional predicate need to be invented.
(3). The learned model has $\alpha\cap I_{\omega}\neq\emptyset,\alpha\nsubseteq I_{\omega},I_{\omega}\nsubseteq\alpha$. Then, for each section that initial states of new transitions can fall in:
$\mathbf{I_{\omega}}$$\mathbf{U}$$s_{1}$$s_{2}$aabbccdd
  • (a): $a\notin\alpha,a\notin I_{\omega}$.
    Thus, $\nexists\>\langle s,\omega,s^{\prime}\rangle\text{ such that }\mathbf{1}_{\alpha}(a)=\mathbf{1}_{\alpha}(s)$ while $I_{\omega}(a)\neq I_{\omega}(s)$. No additional predicate need to be invented.
  • (b): $b\notin\alpha,a\in I_{\omega}$.
    Thus, $\mathbf{1}_{\alpha}(b)=\mathbf{1}_{\alpha}(s_{1})$ while $I_{\omega}(a)\neq I_{\omega}(s)$. New predicate will be invented.
  • (c): $a\in\alpha,a\in I_{\omega}$.
    Thus, $\nexists\>\langle s,\omega,s^{\prime}\rangle\text{ such that }\mathbf{1}_{\alpha}(a)=\mathbf{1}_{\alpha}(s)$ while $I_{\omega}(a)\neq I_{\omega}(s)$. No additional predicate need to be invented.
  • (d): $d\in\alpha,a\notin I_{\omega}$.
    Thus, $\mathbf{1}_{\alpha}(d)=\mathbf{1}_{\alpha}(s_{1})$ while $I_{\omega}(d)\neq I_{\omega}(s)$. New predicate will be invented.
So far, we have discussed all possible cases where initial states of new transitions can fall in, and the predicate invention condition is proven to handle all cases.
Appendix E Additional Details on the Hypothesis Class and Sample Complexity
Recall that our learned symbolic model has the form
$\mathcal{M}=(\mathcal{P},\mathcal{A}),$
where $\mathcal{P}$ is a set of predicates and $\mathcal{A}=\bigcup_{\omega\in\Omega}\mathcal{A}_{\omega}$ is a set of abstract operators, with $\mathcal{A}_{\omega}$ the set of operators associated with skill $\omega\in\Omega$.
Each operator $a\in\mathcal{A}_{\omega}$ is defined by its preconditions and add/delete effects:
$a\equiv\bigl(\omega,\Theta_{a},\textsc{Pre}_{a},\textsc{Eff}_{a}^{+},\textsc{Eff}_{a}^{-}\bigr),\quad\textsc{Pre}_{a},\textsc{Eff}_{a}^{+},\textsc{Eff}_{a}^{-}\subseteq\mathcal{P}.$
In our implementation, the number of VLM calls is finite, and the resulting models use a small number of predicates in environments with finitely many objects. For the theoretical analysis, we make this implicit resource bound explicit:
  • Let $P_{\max}$ denote a fixed maximum number of predicates that SkillWrapper is allowed to invent.
  • Let $\mu_{\max}$ denote the maximum arity of any predicate $\sigma\in\mathcal{P}$.
Because SkillWrapper learns one operator per lifted effect cluster, we can derive an upper bound on the number of operators per skill $\omega\in\Omega$ based on the number of possible effect sets. For any operator $a\in{\mathcal{A}}_{\omega}$ and predicate $\sigma\in{\mathcal{P}}$, there are three possible cases: $\sigma\in\textsc{Eff}_{a}^{+}$, $\sigma\in\textsc{Eff}_{a}^{-}$, or $\sigma\notin\textsc{Eff}_{a}^{-}\cup\textsc{Eff}_{a}^{+}$. Because the upper bound of possible instances of $p$ is $|\mathcal{O}|^{\mu_{\max}}$, we can express the maximum number of operators as
$A_{\max}=3^{P_{\max}\cdot|\mathcal{O}|^{\mu_{\max}}}$
We therefore define the hypothesis class analyzed in Theorem 2 as
$\mathcal{H}=\Bigl\{(\mathcal{P},\mathcal{A})\,\Big|\,|\mathcal{P}|\leq P_{\max},\;|\mathcal{A}_{\omega}|\leq A_{\max}\ \ \forall\,\omega\in\Omega\Bigr\}.$ (7)
Predicate re-evaluation and removal during learning do not expand $\mathcal{H}$; they only move the learned model within this resource-bounded class by altering which predicates and operators are actively used.
E.1 Upper Bound on $|\mathcal{H}|$
We now derive a practical upper bound on the size of $\mathcal{H}$ in equation 7.
Fix a predicate set $\mathcal{P}$ with $|\mathcal{P}|\leq P_{\max}$. For each operator $\alpha$, its symbolic definition is given by three subsets of $\mathcal{P}$:
$\textsc{Pre}_{a},\ \textsc{Eff}_{a}^{+},\ \textsc{Eff}_{a}^{-}\subseteq\mathcal{P}.$
We consider negative precondition, so there are three possibilities for one predicate $p$: $p\in\textsc{Pre}_{a}$, $p\in\textsc{Pre}_{a}$, and $p\notin\textsc{Pre}_{a}$. Since $\textsc{Eff}_{a}^{+}$ and $\textsc{Eff}_{a}^{-}$ are considered in $A_{\max}$, a single operator has at most
$3^{P_{\max}\cdot|\mathcal{O}|^{\mu_{\max}}}$ (8)
distinct configurations of preconditions.
For a fixed skill $\omega\in\Omega$, we allow at most $A_{\max}$ operators. Treating each of the $A_{\max}$ operator “slots” as independently choosing one of the $3^{P_{\max}|\mathcal{O}|^{\mu_{\max}}}$ possible configurations in equation 8, the total number of operator-sets $\mathcal{A}_{\omega}$ for that skill is bounded by
$\bigl(3^{P_{\max}\cdot|\mathcal{O}|^{\mu_{\max}}}\bigr)^{A_{\max}}\;=\;3^{P_{\max}A_{\max}|\mathcal{O}|^{\mu_{\max}}}.$ (9)
Across all skills $\omega\in\Omega$, we obtain the bound
$|\mathcal{H}|\;\leq\;\prod_{\omega\in\Omega}3^{P_{\max}A_{\max}|\mathcal{O}|^{\mu_{\max}}}\;=\;3^{P_{\max}A_{\max}|\Omega||\mathcal{O}|^{\mu_{\max}}}.$ (10)
E.2 Sample Complexity for a Target $(\epsilon,\delta)$
Theorem 2 states that, for any $\epsilon>0$,
$\Pr\Bigl[d_{\mathrm{compl}}\bigl(\widehat{\mathcal{M}}_{n},\mathcal{M}^{*}\bigr)>\epsilon\Bigr]\;\leq\;|\mathcal{H}|\,e^{-n\epsilon},$ (11)
where $\widehat{\mathcal{M}}_{n}$ is the model returned by SkillWrapper after observing $n$ i.i.d. transitions, and $d_{\mathrm{compl}}$ is the completeness distance defined in the main text (probability of a “missed feasible” event under the transition distribution).
Substituting the bound on $|\mathcal{H}|$ from equation 10 into equation 11 yields
$\Pr\Bigl[d_{\mathrm{compl}}\bigl(\widehat{\mathcal{M}}_{n},\mathcal{M}^{*}\bigr)>\epsilon\Bigr]\;\leq\;3^{P_{\max}A_{\max}|\Omega||\mathcal{O}|^{\mu_{\max}}}\,e^{-n\epsilon}.$ (12)
To guarantee that this probability is at most $\delta\in(0,1)$, it suffices that
$3^{P_{\max}A_{\max}|\Omega||\mathcal{O}|^{\mu_{\max}}}\,e^{-n\epsilon}\;\leq\;\delta,$
which is equivalent to
$n\;\geq\;\frac{1}{\epsilon}\Bigl(P_{\max}A_{\max}|\Omega||\mathcal{O}|^{\mu_{\max}}\ln 3+\ln\tfrac{1}{\delta}\Bigr).$ (13)
Thus, the number of transitions required to ensure
$d_{\mathrm{compl}}\bigl(\widehat{\mathcal{M}}_{n},\mathcal{M}^{*}\bigr)\leq\epsilon\quad\text{with probability at least }1-\delta$
is
$n=\mathcal{O}\!\left(\frac{P_{\max}A_{\max}|\Omega||\mathcal{O}|^{\mu_{\max}}+\log(1/\delta)}{\epsilon}\right).$ (14)
In our experiments, the arity and realized numbers of predicates are small. Setting $P_{\max}$ to match the practical budget yields numerical values in equation 13 for the regimes we study. Our theoretical result therefore formalizes how increasing the capacity of the symbolic model (via larger $P_{\max}$) trades off against the number of transitions needed to achieve a desired completeness level.
Appendix F Learned Operators, Case Studies, and Example Tasks
F.1 Learned Operators
# Example Learned Predicate and Operator of Burger domain
# Predicate
name: cut_into_pieces
types:
- cuttable
semantic: "thecuttableappearsasatleasttwonon-touchingvisiblepieces(hasmultipledisconnectedregions),indicatingitiscutratherthanasingleintactpiece."
# Operator
(:action Cut_7
:parameters (?cuttable_p0 - cuttable
?pickupable_p1 - pickupable
?pickupable_p2 - pickupable
?pickupable_p6 - pickupable
?robot_p5 - robot
?station_p4 - station)
:precondition (and
(not (= ?pickupable_p1 ?pickupable_p2))
(not (= ?pickupable_p1 ?pickupable_p6))
(not (= ?pickupable_p2 ?pickupable_p6))
(gripper_empty ?robot_p5)
(on_cutting_board ?cuttable_p0)
(on_station ?cuttable_p0 ?station_p4)
(top_most ?cuttable_p0)
(not (cut_into_pieces ?cuttable_p0))
(not (holding ?robot_p5 ?cuttable_p0))
(not (holding ?robot_p5 ?pickupable_p1))
(not (holding ?robot_p5 ?pickupable_p2))
(not (stacked_on ?pickupable_p6 ?cuttable_p0)))
:effect (and
(cut_into_pieces ?cuttable_p0)))
# Example Learned Predicate and Operator of Franka domain
# Predicate
name: plate_is_dirty
types:
- plate
semantic: the specified plates upper surface contains visible food traces whose appearance differs from the plates base surfac (visible residue to be wiped).
# Operator
(:action Stack_5
:parameters (?pickupable_p0 - pickupable
?pickupable_p1 - pickupable
?pickupable_p3 - pickupable
?plate_p4 - plate
?robot_p2 - robot)
:precondition (and
(not (= ?pickupable_p0 ?pickupable_p1))
(not (= ?pickupable_p0 ?pickupable_p3))
(not (= ?pickupable_p1 ?pickupable_p3))
(holding ?robot_p2 ?pickupable_p1)
(plate_top_unoccupied ?plate_p4)
(not (gripper_empty ?robot_p2))
(not (holding ?robot_p2 ?pickupable_p0))
(not (holding ?robot_p2 ?pickupable_p3))
(not (plate_is_dirty ?plate_p4))
(not (stacked_on ?pickupable_p1 ?plate_p4)))
:effect (and
(gripper_empty ?robot_p2)
(stacked_on ?pickupable_p1 ?plate_p4)
(not (holding ?robot_p2 ?pickupable_p1))
(not (plate_top_unoccupied ?plate_p4))))
# Example Learned Predicate and Operator of Bi-manual Kuka domain
# Predicate
name: Coated
types:
- utensil
semantic: a visible layer or clump of material adheres to the utensils working end (e.g., the blade shows a smear that was absent before).
# Operator
(:action Scoop_156
:parameters (?openable_p0 - openable
?robot_p1 - robot
?utensil_p2 - utensil)
:precondition (and
(HeldByRobot ?robot_p1 ?openable_p0)
(InLeftGripper ?robot_p1 ?openable_p0)
(InRightGripper ?robot_p1 ?utensil_p2)
(LidOff ?openable_p0)
(not (Closed ?openable_p0))
(not (Coated ?utensil_p2))
(not (InContainer ?utensil_p2))
(not (LeftGripperEmpty ?robot_p1))
(not (OpenableOnTable ?openable_p0))
(not (RightGripperEmpty ?robot_p1))
(not (UtensilOnTable ?utensil_p2)))
:effect (and
(Coated ?utensil_p2)))
Additionally, we present here a predicate and an operator written by the PDDL expert:
# Predicate and Operator Written by PDDL Expert
# Predicate
name: is_on_station
types:
- pickupable
- station
semantic: "A‘pickupable‘objectisontopofa‘station‘."
# Operator
(:action Cut
:parameters (?robot - robot
?cuttable - cuttable
?board - cuttingboard)
:precondition (and
(hand_empty)
(obj_free ?cuttable)
(is_on_station ?cuttable ?board)
(not (is_cut ?cuttable)))
:effect (and
(is_cut ?cuttable)))
F.2 Result Analysis of Robotouille
We here present case studies of the failure modes of SkillWrapper and the baselines. One common failure mode is misclassifications during inference time, where we abstract low-level states specified in raw images into symbolic states using the predicates in the learned model. Since our contributions are more on the theory and algorithm design, we did not explore various potential techniques, such as ViperGPT (Surís et al., 2023), that can possibly mitigate this issue.
System Predicates (Sys Preds.) are designed in a way such that every possible state can be specified with a compact set of eight predicates, as shown in the experiments. However, this is proven insufficient to describe the model dynamics. For example, the predicate $\mathtt{clear\_above}$ or its semantic equivalent is necessary for describing the precondition of all operators learned from the Pick skill, such that the model can express “Picking an item with another item on top will fail.” Since predicate invention is disabled in Sys Preds., it often returns invalid plans that attempt to pick objects that are underneath a stack. As a result, the lack of soundness induces a low solved rate for hard and impossible problems of System Predicates.
Random exploration is heavily limited by inefficient data gathering. However, if we take a closer look at the operators it learns, they are of good quality, which is the benefit of sharing the same predicate invention algorithm of SkillWrapper. In fact, one frequent failure mode under this randomness is that the baseline may never execute the skills with complex preconditions successfully. For example, in the Robotouille setting, the Cook action requires the item to be on the stove and the agent’s hand to be empty. Since the same predicate invention algorithm with SkillWrapper is being used, it can only learn operators for the skills that have been successfully executed in the observed transitions. As a result, it only learns operators for skills that are usually executable, such as Pick, and it thus can only solve the simplest pick and stack tasks. Additionally, random exploration also achieves $100\%$ on Impossible, yet because the learned model is not plannable.
No Heuristic shares a similar failure mode as System Predicates—achieves good performance in Easy problems but degrades significantly in Hard. With further investigation, we found that the baseline generally invents fewer predicates than SkillWrapper, which results in occasionally missing critical ones (two out of the five total runs). This observation explains the large variances in Hard and Impossible problems, and also indicates that Easy problems could be solved even with an incomplete predicate set, which aligns with our findings in System Predicate. In turn, it supports the usefulness of the two engineered heuristics for skill sequence proposal. We believe improving the exploration strategy is a promising direction for future work. Another point we want to note is that either the solved rate or the planning budget only evaluates the learned operators from the planning outcome, while better metrics are needed for evaluating the exploration efficiency.
SkillWrapper’s failure mode is similar to the case of Random exploration. Comparing the poorly performing PDDL operators learned by SkillWrapper to Expert Operators, the model only contains one extra predicate, $\mathtt{on\_cutting\_board(item)}$, which divides the previous cluster that shares the same effect into even smaller clusters. Then, the transition data in these smaller clusters cannot support the model learning algorithm to effectively eliminate spurious preconditions. This case study points out that the balance between predicate invention and data gathering is a critical factor in the learning process: if too many predicates are invented without adequate transition data, the resulting operators could possibly contain spurious preconditions, such that it cannot generalize at all. In SkillWrapper, the balance is controlled by empirically tuning the threshold of the scoring function and the length of each skill sequence proposed. Another sub-optimality is that our algorithm does not filter invented predicates that are semantic synonyms or antonyms to existing predicates, which increases the classification burden of the foundation model. Though these redundant predicates are usually well handled by the foundation model, a smarter prompting system could be designed to mitigate this issue and improve computation efficiency.
F.3 Example Tasks
Here, we provide additional examples of the tasks, using planning problem from the evaluation data. Specifically, Robotouille tasks use images generated by the simulator, Franka tasks use images taken by a fixed camera in front of the robot, and Bimanual Kuka tasks use images taken by its egocentric camera.
(a) Initial state
(b) Goal state
Figure 7: Example task in Robotouille.
(a) Initial state
(b) Goal state
Figure 8: Example task in Franka.
(a) Initial state
(b) Goal state
Figure 9: Example task in Bimanual Kuka.
Appendix G VLM Reliability Study
G.1 Classification Accuracy
We here analyze the classification accuracy of the vision-language model (VLM) used in both robotic experiments. Since the predicates are generated on the fly by the VLM, we do not have the ground truth values for them, and thus we must verify if the truth values match the images manually.
Per-predicate Classification Evaluation.
Since predicates are originally lifted and can be grounded with different combinations of objects, we first define a classification over a low-level state of a grounded predicate as correct if (1) all parameters appear in the scene (if the predicate is not nullary) and (2) the truth value of the predicate match the low-level state specified by the image input. Then, we define a classification of a lifted predicate over a low-level state as correct if all of its grounded instances are classified correctly over that state.
Results and Analysis.
Over all predicates, the classification accuracy is $86.7\%$ for the Franka experiment, and $98.5\%$ for the bimanual Kuka experiment. Compared to the planning performance reported for both experiments, the classification accuracy is generally much higher. One reason for this mismatch is that, due to the rigidity of symbolic planning, even flipping the truth value of a single predicate can lead to a planning failure. To support this claim, we found specific poorly performing predicates that hinder the planning task the most, and we provide more quantitative results in the next paragraph.
Per-predicate Accuracy.
The learned symbolic model of the Franka experiment contains $6$ predicates, which have $11$ possible grounded instances. The learned symbolic model of the bimanual Kuka experiment contains $12$ predicates, which have $13$ possible grounded instances. We evaluate per-predicate accuracy for both in Table 3 and Table 4. From the results of the Franka experiment, we identify the two predicates, $\mathtt{gripper\_empty}$ and $\mathtt{holding}$, that caused all planning failures, and they fail almost simultaneously due to their semantic correlation. With further investigation, we found that the misclassifications were induced by a single object, $\mathtt{Sponge}$, which is possibly due to the color of the object and the background being too similar. In the bimanual Kuka experiment, it is $\mathtt{coated}$ (if the knife has peanut butter on it) that caused most of the planning failure, likely caused by the lighting conditions. These observations suggest the accuracy of VLM is a limiting factor, and resolving them poses a promising path to improving the performance of SkillWrapper.
Table 3: Per-predicate accuracy of Franka.
gripper_empty holding mug_full plate_top_unoccupied stacked_on plate_is_dirty
Accuracy (%) 60.0 60.0 100.0 100.0 100.0 100.0
Table 4: Per-predicate accuracy of Bimanual Kuka.
InLeftGripper InRightGripper RightGripperEmpty LeftGripperEmpty LidOff InContainer
Accuracy (%) 100.0 100.0 100.0 100.0 100.0 98.3
OpenableOnTable Closed Coated SpreadOn HeldByRobot UtensilOnTable
Accuracy (%) 96.7 100.0 88.3 98.3 100.0 100.0
(a) $\mathtt{holding(Sponge)}=\mathtt{\textbf{F}}$
(b) $\mathtt{coated(Knife)}=\mathtt{\textbf{T}}$
G.2 Real-world Robustness
To evaluate the real-world robustness of the VLM, we additionally conduct experiments to investigate factors such as viewpoints, lighting conditions, or domain shifts. For each of them, we collect a held-out set of images by varying these factors. We report per-predicate accuracy, and all numbers are averaged across three individual runs.
[…]

### v4

@@ III-D Theoretical Analysis
Sketch. Since $\widehat{\mathrm{Err}}(\widehat{{\mathcal{M}}}_{n})=0$ by construction, a Chernoff bound and union bound over the finite hypothesis class imply that the true error is small with high probability. (We derive the upper bound of the hypothesis class in Appendix -E.)
IV Experiments
(a) Franka: Initial
(b) Franka: Goal
(c) Bimanual Kuka: Initial
(d) Bimanual Kuka: Goal
Figure 3: Initial and goal states for robot experiments.
IV-A Implementation of Predicates
  • A foundation model gives us a string that can be used as a lifted predicate (in other words, it generates interpretable relational classifiers with a good heuristic).
With these two properties, we can use the VLM’s response in string form as a relational predicate, and a grounded version of the predicate can be used as a classifier. We use GPT-5 [47] for predicate generation and evaluation. In addition to the system performance reported in the main paper, we also conducted comprehensive studies of the component-wise reliability of the VLM in Appendix -I, including predicate invention and classification tests, per-predicate accuracy, and robustness to lighting conditions and view angle changes. This proves the real-world robustness of SkillWrapper.
(a) Robotouille: Initial
(b) Robotouille: Goal
Figure 4: Initial and goal state images in the Robotouille domain.
IV-B Simulation
We first conduct experiments in Robotouille [19], a simulated grid-world kitchen domain with a robot that has five high-level skills: Pick, Place, Cut, Cook, and Stack. In the environment, there are several objects: a patty, lettuce, a top bun, and a bottom bun; there is also a cutting board and a stove for cutting the lettuce and cooking the patty, respectively. We design and categorize 50 abstract planning problems: 20 easy problems, whose optimal solutions have no more than 7 steps; 20 hard problems, whose optimal solutions have no more than 15 steps; and 10 impossible tasks that cannot be realized in the environment, which a sound model must identify.
For a complete evaluation of SkillWrapper, we compare SkillWrapper against the following baselines:
  • Expert Operators: A human expert familiar with PDDL is asked to interact with the environment and manually write predicates with semantics and PDDL operators.
  • System Predicates: This baseline directly uses the built-in predicate set of the simulator, which is designed to define any possible simulated state unambiguously. In addition, instead of evaluating the predicate truth values with foundation models, this baseline has access to the exact abstract states of the simulator. However, predicate invention is disabled in this environment.
  • ViLa [24]: This baseline is a closed-loop VLM-based approach that iteratively prompts a foundation model for the next action until the goal state is reached, given an image observation and the robot’s action history.
  • Random Exploration: Instead of proposing skill sequences, this baseline, acting as a lower-bound, randomly samples a skill and populates the arguments with valid objects. This baseline shares the same predicate invention and operator learning algorithms as SkillWrapper.
  • FM Invent: This baseline invents predicates in batches by directly prompting the foundation model with all image states, instead of using the predicate invention method of SkillWrapper. The active data collection and operator learning procedures are the same as SkillWrapper.
  • No Heuristic: This baseline is the same as SkillWrapper, except that skill sequences are selected randomly from the foundation model’s output without applying the heuristics (Sec. III-A).
For each baseline that performs operator learning, we run the learning algorithm for five iterations, with each iteration proposing and executing one skill sequence comprised of 15 steps as its interaction budget. We then evaluate each method on the evaluation set and report the average results in Table I, where Solved % is the percentage of the problem set that was successfully solved or where impossible tasks were correctly identified by returning an empty plan, and PB stands for planning budgets—the number of plans that were tried before solving the problem (adopted from [39]). We set a planning budget cap of 10 across all problems; if the planning budget has been used up for a problem, we consider it a failure. Theoretically, PB is an adequate metric that reflects the completeness of the learned model, and the impossible problems reflect its soundness. As shown in the table, SkillWrapper outperforms all baselines that have no access to privileged knowledge, and even surpasses the performance of the System Predicates baseline on hard problems. For brevity, we leave case studies, example operators, failure modes, and analysis of simulated results in Appendix -H.
TABLE I: Baseline Comparison in Robotouille Environment
Method Easy Hard Impossible Solved % $\uparrow$ PB $\downarrow$ Solved % $\uparrow$ PB $\downarrow$ Expert Ops. 81.0 $\pm$ 3.7 1.9 $\pm$ 0.4 58.1 $\pm$ 3.9 4.2 $\pm$ 0.4 100 $\pm$ 0.0 Sys Preds. 79.0 $\pm$ 3.7 2.6 $\pm$ 0.2 22.0 $\pm$ 12.9 7.8 $\pm$ 1.3 42.0 $\pm$ 7.5 ViLa [24] 46.0 $\pm$ 16.2 - 13.9 $\pm$ 11.6 - 20.0 $\pm$ 10.9 Random Exp. 4.0 $\pm$ 2.0 9.6 $\pm$ 0.2 0 $\pm$ 0.0 10.0 $\pm$ 0.0 100 $\pm$ 0.0 FM Invent 23.3 $\pm$ 16.5 7.7 $\pm$ 1.7 1.7 $\pm$ 2.4 9.7 $\pm$ 0.2 63.3 $\pm$ 26.2 No Heuristic 76.0 $\pm$ 4.9 2.5 $\pm$ 0.9 24.0 $\pm$ 19.6 7.8 $\pm$ 1.8 80 $\pm$ 20.9 Ours 74.0 $\pm$ 3.7 2.7 $\pm$ 0.4 40.0 $\pm$ 3.2 6.3 $\pm$ 0.4 100 $\pm$ 0.0
IV-C Real Robots
To demonstrate the applicability of SkillWrapper for real-world robotic settings, we designed two sets of experiments with two robotic platforms: a Franka Emika Panda robot (Figures 3(a) and 3(b)) and a bimanual platform with two Kuka iiwa robots (Figures 3(c) and 3(d)). For both robot experiments, we assume that all skills are deterministic, which is a common assumption made by existing work [57, 22, 39, 4] (see Appendix -F for further details). We used an interaction budget of 15 steps per iteration for all methods requiring active data collection. The supplementary material contains video demonstrations of both experiments.
Generalization of SkillWrapper
In this setting, a tabletop Franka Emika Research 3 (Panda) manipulator has a skill set $\Omega$ consisting of five black-box skills: Pick, Place, Stack, Pour, and Wipe. The object set $\mathcal{O}$ contains five objects: a mug, a teapot, a plate, a bowl, and a sponge. The robot can pick up and place all objects except the mug and plate, pour ingredients from the teapot into the mug, and use the sponge to wipe the plate if it is dirty. To validate the generalization ability of the learned skill representations, we design three smaller training environments, where each environment contains a subset of $\mathcal{O}$, and thus only a subset of $\Omega$ is executable. Each training environment contains fewer than 10 possible states. After running SkillWrapper for exactly one iteration for each environment, we port the learned skill representations to the test environment that contains all objects in $\mathcal{O}$, which induces a state space of 34 possible abstract states. To quantify this generalization process, we similarly prepare an evaluation problem set that consists of five problems in the training environments, five problems in the test environment, and five Impossible problems across the two environments. The results are shown in Table II.
TABLE II: Results of Generalization Experiment
[…]